1. Determine the next term of the following sequence: 0, 3,8,15...

1.1. What is the eleventh term of this sequence? Show all your working.

1.2. Name the type of numbers formed by the numbers added in the sequence.
1
Expert's answer
2020-07-09T19:35:14-0400

Let "a_n" be the "n" -th element of the sequence: "0,3,8,15,\u2026"

Then "a_n=n^2-1"

"a_1=1^2-1=0"

"a_2=2^2-1=3"

"a_3=3^2-1=8"

"a_4=4^2-1=15"

1.The next term is "a_5=5^2-1=24" ,

1.1 The eleventh term of this sequence is "a_{11}=11^2-1=121-1=120" , "a_{11}=120"

1.2 "a_{n+1}-a_n=((n+1)^2-1)-(n^2-1)=n^2+2n+1-1-n^2+1=2n+1"

Type of numbers: odd numbers.